# Power-law banded random matrix ensemble as a model for quantum many-body Hamiltonians

### v2

@@ Abstract
We explore interpretations of the power-law banded random matrix (PLBRM) ensemble as Hamiltonians of one-dimensional quantum many-body systems. We introduce and compare a number of labeling schemes for assigning random matrix basis indices to many-body basis vectors. We compare the physical properties of the resulting Hamiltonians, focusing on the half-system eigenstate bipartite entanglement entropy. We show and quantify how the different PLBRM phases (ergodic, weakly ergodic, localized), known from the single-particle interpretation, can be interpreted as entanglement transitions in the quantum many-body interpretation. For the weakly ergodic phase, where spectral edge and bulk eigenstates show distinct behavior, we perform a detailed scaling analysis to provide a quantitative picture of the boundaries between different types of entanglement scaling behaviors. In particular, we identify and characterize an intermediate set of eigenstates whose entanglement entropy have volume law scaling but non-vanishing deviation from the Page value expected for maximally ergodic states.
@@ I Introduction
Ideas from random matrix theory [1, 2, 3] are central to our current understanding of isolated quantum systems and quantum many-body chaos [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Away from the edges of the many-body spectrum, important aspects of physical eigenstates and eigenvalues are well-described by random matrix models, provided the system is not integrable or many-body localized.
The most prominent random matrix ensembles used to model quantum Hamiltonians are the Gaussian orthogonal (GOE) and unitary (GUE) ensembles, for which the matrix elements are uncorrelated and identically distributed. However, in many physical situations, remote basis states are on average less strongly coupled than nearby basis states. Several random matrix models or their variants incorporate this idea, e.g., the Rosenzweig-Porter model [15, 16, 17, 18, 19, 20, 21] and recently proposed variations thereof [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], the $\beta$ ensemble [33, 34, 35, 36, 37, 38], banded random matrices [39, 40, 41, 42, 43, 44, 45, 46, 47], ultrametric random matrix models [48, 49, 50, 51, 52, 53, 54, 55, 56], and power-law banded random matrix (PLBRM) models [57, 58, 59, 60, 61, 62, 49, 50, 52, 54, 63, 64, 65, 66, 67, 68, 69, 70, 71]. The last is the focus of the present work.
@@
In this work we explore many-body interpretations of the power-law banded random matrix ensemble. The usual random-matrix description of many-body Hamiltonians with matrix classes GOE or GUE has the shortcoming that the eigenvectors of such random matrices have similar properties in the spectral bulk and at the spectral edges. In contrast, for physical many-body Hamiltonians, the low-energy eigenstates have markedly different properties compared to the eigenstates in the spectral bulk. Perhaps the best known manifestation of this edge-bulk distinction is seen in the entanglement entropy ($S_{\text{ent}}$) of the eigenstates: At the spectral edges, $S_{\text{ent}}$ is low (‘area law’) [80, 81], while in the mid-spectrum ‘infinite-temperature’ regime, the eigenstates have $S_{\text{ent}}$ close to the value expected for random states, the so-called Page value $S_{\text{Page}}$ (‘volume law’). As a result, for chaotic many-body systems, a scatter plot of $S_{\text{ent}}$ versus eigenenergy takes the shape of an arch or rainbow, by now familiar from many numerical examples [82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 67, 95, 96, 97, 98, 99]. The absence of this structure in GOE or GUE random matrices is rather striking, and is a motivation to study the many-body interpretation of other, more structured, random matrix ensembles. In particular, it has been noted [67] that power-law banded random matrices reflect this distinction between edge and bulk eigenstates in the weakly ergodic phase. Thus, this random matrix ensemble provides an improved model of the structure of chaotic many-body Hamiltonians. This background also provides motivation to study in detail the entanglement entropy of the eigenstates of PLBRMs with a many-body interpretation. This is a main focus of the present work.
@@
In the present work, we provide a more systematic study of interpreting PLBRMs as Hamiltonians of quantum many-body systems. After a brief review of PLBRMs (Sec. II), we first identify three different ways to label basis vectors with many-body configurations (Sec. III). We quantify the quality of the labeling schemes, and argue one of our labeling schemes to improve on the previously employed binary scheme. We show how to resolve an issue regarding the homogeneity of the resulting Hamiltonians using what we will refer to as ‘site randomization’. Next (Sec. IV), we turn to the physical properties of the resulting Hamiltonians, focusing on the entanglement entropy. We study the scaling of the eigenstate entanglement entropy with system size in the different phases of the model.
The general picture is that $S_{\text{ent}}$ exhibits volume law behavior throughout the spectrum in the fully ergodic phase ($\alpha<\frac{1}{2}$), area law behavior throughout the spectrum in the localized phase ($\alpha>1$), while in the intermediate weakly ergodic phase ($\frac{1}{2}<\alpha<1$) there is a bulk-edge difference with the bulk showing volume law and the edge showing area law. We confirm this picture quantitatively through separate finite-size scaling analyses for bulk $S_{\text{ent}}$ and edge $S_{\text{ent}}$ (Sec. V). The weakly ergodic phase shows the most interesting variation between the different parts of the spectrum, and this is also the behavior familiar from many-body systems. Therefore we undertake a more detailed study of this phase, seeking to identify the demarcation between bulk and edge eigenstates (Sec. VI). Using scaling analysis on different quantities, we numerically identify an intermediate set of eigenstates, which have volume law scaling but a non-vanishing deviation from the Page value. There are thus two boundaries in energy. One boundary separates eigenstates whose entanglement entropy approaches the Page value ($S_{\text{ent}}{\to}S_{\text{Page}}$) from these intermediate eigenstates. A second boundary separates these intermediate eigenstates from the area law eigenstates at the very edge. We conclude (Sec. VII) with a summary and discussion and some questions opened up by this work for future investigations.
@@ II Power-law banded random matrices
where $G_{ij}=G_{ji}$ are independent and identically distributed random numbers, representing the Gaussian orthogonal (GOE) random matrix ensemble. The GOE consists of real-valued symmetric matrices with entries sampled independently from the Gaussian distribution with mean $\mu=0$ and off-diagonal (diagonal) components with variance $\sigma^{2}=1/2$ ($\sigma^{2}=1$[1]. The function $a(r)$ decays as a power law for $r\gg 1$, with a tunable exponent $\alpha>0$. We will use
$a(r)=\frac{1}{1+(r/\beta)^{\alpha}}$ (2)
with the so-called bandwidth $\beta$ set to unity ($\beta=1$). This definition has been part of investigations previously in Ref. [67], involving some of the present authors. Other often used definitions can be found, for example, in Refs. [57, 58, 60, 59, 61, 49, 54, 68, 69]. A few of these alternative forms are of the type $a(r)\sim 1/\sqrt{1+(r/\beta)^{2\alpha}}$ or $a(r)\sim 1/\left[1+(r/\beta)^{2}\right]^{\alpha/2}$, or periodic versions thereof. Like the form (2) that we are using, these are also smooth functions satisfying the same limiting behaviors: $r\sim\text{const}$ for $r\ll b$ and $r\sim r^{-\alpha}$ for $r\gg b$. We expect our qualitative results to be independent of the explicit choice of $a(r)$, as long as the bandwidth $\beta$ is not large.
In the single-particle interpretation, the parameter $\alpha$ allows one to tune the eigenstates in the bulk (middle) of the spectrum in a fully ergodic ($\alpha<\frac{1}{2}$[53], weakly ergodic ($\frac{1}{2}<\alpha<1$[54], or (power-law) localized ($\alpha>1$) phase [57]. This $\alpha$-dependence is illustrated in Fig. 1. In the fully ergodic phase, the properties of the model are statistically the same as those of the GOE. This means that for an eigenstate $|\psi\rangle$, the inverse participation ratio $\text{IPR}=\sum_{n}|\langle n|\psi\rangle|^{4}$, with the summation running over all basis states $|n\rangle$, is asymptotically given by $3/N$, where $N$ is the matrix dimension ($n=1,2,\dots,N$). The weakly ergodic phase is characterized by bulk eigenstates occupying only a finite fraction of the Hilbert space [54, 100]. Here, the inverse participation ratio of the eigenstates is asymptotically given by $c/N$ for some $c>3$ depending on $\alpha$. The fractal dimensions $D_{q}$ ($q\geq 1$) of an eigenstate $|\psi\rangle$ are defined through the scaling
@@
The fully and weakly ergodic phases are characterized by eigenstates with unit fractal dimensions ($D_{q}=1$). At the critical point $\alpha=1$, the eigenstates in the bulk of the spectrum are multifractal ($10<D_{q}<$ 1 is a $q$-dependent function), and level statistics are intermediate between Poisson and Wigner-Dyson. In the localized phase, the eigenstates have fractal dimension zero for positive integers $q$, meaning that the inverse participation ratio does not scale as $1/N$ and the eigenstates only occupy a measure zero fraction of the full Hilbert space. Fig. 1 provides a graphical summary of the phase diagram discussed here.
Figure 1: The phase diagram of the power-law banded random matrix ensemble, showing the phases of mid-spectrum eigenstates as a function of the power-law exponent $\alpha$. The panels show the eigenstate entanglement entropy for a single realization as a function of the normalized eigenstate index (random labeling scheme, $L=10$), for the three phases ($\alpha=0.25$, $\alpha=0.75$, $\alpha=1.25$). The horizontal dashed line indicates the Page entanglement entropy $S_{\text{Page}}$. (Ensemble-averaged versions of such plots are discussed later, see Fig. 4.)
The phenomenology of power-law banded random matrices can be understood from resonance counting [101, 102] and the Breit-Wigner approximation (see, for example, Refs. [103, 21]). First, from resonance counting arguments, it follows that localization perturbation theory converges if the number $N_{\text{res}}$ of resonances, defined as $|H_{mn}|>|H_{nn}-H_{mm}|$, does not grow with the size $N$ of the matrix [23, 25, 64]. To calculate $N_{\text{res}}$, one should sum the probabilities of the other nodes $m\neq n$ to be in resonance with node $n$. For identically, independently uniformly distributed diagonal elements $|H_{ii}|<W/2$, one obtains
[…]
@@
$\Gamma_{n}(E)=\frac{2\pi}{\hbar}\nu(E)\sum_{\begin{subarray}{c}m=1\\ m\neq m\end{subarray}}^{N}|H_{mn}|^{2},$ (5)
@@
with a certain normalization prefactor $A$ [103, 21]. As soon as $\Gamma(E)\equiv\left\langle\Gamma_{n}(E)\right\rangle\gg\left\langle H_{nn}^{2}\right\rangle$ for all the energies $E$, all the wave-function coefficients are of the same order and, therefore, all the states should be ergodic. For power-law banded random matrices, one obtains $\Gamma_{n}(E)\sim\sum_{R=1}^{N}R^{-2\alpha}\sim N^{1-2\alpha}+\mathcal{O}(1)$. This diverges at $\alpha<\frac{1}{2}$ for all the energies. Therefore, for $\alpha<\frac{1}{2}$, one expect all the eigenstates to be ergodic (statistically equivalent to those of the GOE). Together, these two lines of arguments predict the ‘phase diagram’ shown in Fig. 1.
@@
In the quantum many-body interpretation, to be developed in the following sections, the two transitions appear to correspond to entanglement phase transitions. As one increases $\alpha$, we start off with all eigenstates having volume law behavior for $\alpha<\frac{1}{2}$. As $\alpha$ crosses $\alpha=\frac{1}{2}$, the edge eigenstates change in nature and acquire area law scaling, and then when $\alpha$ crosses $\alpha=1$, the bulk eigenstates also turn to area law scaling. These notions will be confirmed through numerical scaling analyses later on, but we have provided a visualization of the different behaviors in the three panels in Fig. 1. The panels show single-realization scatter plots of $S_{\text{ent}}$ against eigenvalue index, using what we call the random labeling scheme (Sec. III).

### v3

@@ Abstract
We explore interpretations of the power-law banded random matrix (PLBRM) ensemble as Hamiltonians of one-dimensional quantum many-body systems. We introduce and compare a number of labeling schemes for assigning random matrix basis indices to many-body basis vectors. We compare the physical properties of the resulting Hamiltonians, focusing on the half-system eigenstate bipartite entanglement entropy. We show and quantify how the different PLBRM phases (ergodic, weakly ergodic, localized), known from the single-particle interpretation, can be interpreted as entanglement transitions in the quantum many-body interpretation. For the weakly ergodic phase, where spectral edge and bulk eigenstates show distinct behavior, we perform a detailed scaling analysis to provide a quantitative picture of the boundaries between different types of entanglement scaling behaviors. In particular, we identify and characterize an intermediate set of eigenstates whose entanglement entropy have volume law scaling but nonvanishing deviation from the Page value expected for maximally ergodic states.
@@ I Introduction
Ideas from random matrix theory [1, 2, 3] are central to our current understanding of isolated quantum systems and quantum many-body chaos [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Away from the edges of the many-body spectrum, important aspects of physical eigenstates and eigenvalues are well-described by random matrix models, provided the system is not integrable or many-body localized. The most prominent random matrix ensembles used to model quantum Hamiltonians are the Gaussian orthogonal (GOE) and unitary (GUE) ensembles, for which the matrix elements are uncorrelated and identically distributed.
However, physical Hamiltonians expressed in a natural local basis often exhibit a structure: basis states $|n\rangle$ and $|m\rangle$ tend to be stonger (weaker) coupled when |nm||n-m| is small (large). Several random matrix models or their variants incorporate this idea, e.g., the Rosenzweig-Porter model [15, 16, 17, 18, 19, 20, 21] and recently proposed variations thereof [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], the $\beta$ ensemble [33, 34, 35, 36, 37, 38], banded random matrices [39, 40, 41, 4, 42, 43, 44, 45, 46, 47, 12, 48], ultrametric random matrix models [49, 50, 51, 52, 53, 54, 55, 56, 57], and power-law banded random matrix (PLBRM) models [58, 59, 60, 61, 62, 63, 50, 51, 53, 55, 64, 65, 66, 67, 68, 69, 70, 71, 72]. The last is the focus of the present work.
@@
In this work we explore many-body interpretations of the power-law banded random matrix ensemble. e interpret PLBRMs as Hamiltonians of spin-$\frac{1}{2}$ systems not possessing any conservation laws. This means that, for an $L$-site system, the matrix dimension $N$ is $2^{L}$. The usual random-matrix description of many-body Hamiltonians with matrix classes GOE or GUE has the shortcoming that the eigenvectors of such random matrices have similar properties in the spectral bulk and at the spectral edges. In contrast, for physical many-body Hamiltonians, the low-energy eigenstates have markedly different properties compared to the eigenstates in the spectral bulk. Perhaps the best known manifestation of this spectral edge-bulk distinction is seen in the entanglement entropy ($S_{\text{ent}}$) of the eigenstates: At the spectral edges, $S_{\text{ent}}$ is low (“area law”) [81, 82], while in the mid-spectrum “infinite-temperature” regime, the eigenstates have $S_{\text{ent}}$ close to the value expected for random states, the so-called Page value $S_{\text{Page}}$ (“volume law”). As a result, for chaotic many-body systems, a scatter plot of $S_{\text{ent}}$ versus eigenenergy takes the shape of an arch or rainbow, by now familiar from many numerical examples [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 68, 96, 97, 98, 99, 100]. The absence of this structure in GOE or GUE random matrices is rather striking, and is a motivation to study the many-body interpretation of other, more structured, random matrix ensembles. In particular, it has been noted [68] that power-law banded random matrices reflect this distinction between edge and bulk eigenstates in the weakly ergodic phase. Thus, this random matrix ensemble provides an improved model of the structure of chaotic many-body Hamiltonians. This background also provides motivation to study in detail the entanglement entropy of the eigenstates of PLBRMs with a many-body interpretation. This is a main focus of the present work.
@@
In the present work, we provide a more systematic study of interpreting PLBRMs as Hamiltonians of quantum many-body systems. After a brief review of PLBRMs (Sec. II), we first identify three different ways to label basis vectors with many-body configurations (Sec. III). We quantify the quality of the labeling schemes, and argue one of our labeling schemes to improve on the previously employed binary scheme. We show how to resolve an issue regarding the homogeneity of the resulting Hamiltonians using what we will refer to as “site randomization”. Next (Sec. IV), we turn to the physical properties of the resulting Hamiltonians, focusing on the entanglement entropy. We study the scaling of the eigenstate entanglement entropy with system size in the different phases of the model.
The general picture is that $S_{\text{ent}}$ exhibits volume law behavior throughout the spectrum in the fully ergodic phase ($\alpha<\frac{1}{2}$), area law behavior throughout the spectrum in the localized phase ($\alpha>1$), while in the intermediate weakly ergodic phase ($\frac{1}{2}<\alpha<1$) there is a bulk-edge difference with the bulk showing volume law and the edge showing area law. We confirm this picture quantitatively through separate finite-size scaling analyses for bulk $S_{\text{ent}}$ and edge $S_{\text{ent}}$ (Sec. V). The weakly ergodic phase shows the most interesting variation between the different parts of the spectrum, and this is also the behavior familiar from many-body systems. Therefore we undertake a more detailed study of this phase, seeking to identify the demarcation between bulk and edge eigenstates (Sec. VI). Using scaling analysis on different quantities, we numerically identify an intermediate set of eigenstates, which have volume law scaling but a nonvanishing deviation from the Page value. There are thus two boundaries in energy. One boundary separates eigenstates whose entanglement entropy approaches the Page value ($S_{\text{ent}}{\to}S_{\text{Page}}$) from these intermediate eigenstates. A second boundary separates these intermediate eigenstates from the area law eigenstates at the very spectral edge. We conclude (Sec. VII) with a summary and discussion and some questions opened up by this work for future investigations.
@@ II Power-law banded random matrices
where $G_{ij}=G_{ji}$ are independent and identically distributed random numbers, representing the GOE random matrix ensemble. The GOE consists of real-valued symmetric matrices with entries sampled independently from the Gaussian distribution with mean $\mu=0$ and off-diagonal (diagonal) components with variance $\sigma^{2}=1/2$ ($\sigma^{2}=1$[1]. The function $a(r)$ decays as a power law for $r\gg 1$, with a tunable exponent $\alpha>0$. We will use
$a(r)=\frac{1}{1+(r/\beta)^{\alpha}},$ (2)
with the so-called bandwidth $\beta$ set to unity ($\beta=1$). This definition has been part of investigations previously in Ref. [68], involving some of the present authors. Other often used definitions can be found, for example, in Refs. [58, 59, 61, 60, 62, 50, 55, 69, 70]. A few of these alternative forms are of the type $a(r)\sim 1/\sqrt{1+(r/\beta)^{2\alpha}}$ or $a(r)\sim 1/\left[1+(r/\beta)^{2}\right]^{\alpha/2}$, or periodic versions thereof. Like the form (2) that we are using, these are also smooth functions satisfying the same limiting behaviors: $r\sim\text{const}$ for $r\ll\beta$ and $r\sim r^{-\alpha}$ for $r\gg\beta$. We expect our qualitative results to be independent of the explicit choice of $a(r)$, as long as the bandwidth $\beta$ is not large.
The parameter $\alpha$ can be used to tune the eigenstates in the bulk (middle) of the spectrum. Depending on their behavior, three phases of the system have been identified: the fully ergodic ($\alpha<\frac{1}{2}$[54] phase, the weakly ergodic ($\frac{1}{2}<\alpha<1$[55] phase, and the (power-law) localized ($\alpha>1$) phase [58]. This $\alpha$-dependence is illustrated in Fig. 1.
In the fully ergodic phase ($\alpha<\frac{1}{2}$), the properties of the model are statistically the same as those of the GOE. This means that for an eigenstate $|\psi\rangle$, the inverse participation ratio $\text{IPR}=\sum_{n}|\langle n|\psi\rangle|^{4}$, with the summation running over all basis states $|n\rangle$, is asymptotically given by $3/N$, as is the case with fully ergodic states [1]. Here $N$ is the matrix dimension ($n=1,2,\dots,N$).
The weakly ergodic phase, $\alpha\in\left(\frac{1}{2},1\right)$, is characterized by bulk eigenstates occupying only a finite fraction of the Hilbert space [55, 101]. The inverse participation ratio of the eigenstates is asymptotically given by $c/N$ for some $c>3$ depending on $\alpha$.
The fractal dimensions $D_{q}$ ($q\geq 1$) of an eigenstate $|\psi\rangle$ are defined through the scaling
@@
The fully and weakly ergodic phases are characterized by eigenstates with unit fractal dimensions ($D_{q}=1$). At the critical point $\alpha=1$, the eigenstates in the bulk of the spectrum are multifractal, i.e., $D_{q}\in(0,1)$ is a $q$-dependent function, and level statistics are intermediate between Poisson and Wigner-Dyson.
In the localized phase ($\alpha>1$), the eigenstates have fractal dimension zero for positive integers $q$, meaning that the inverse participation ratio does not scale as $1/N$ and the eigenstates only occupy a finite number of Hilbert-space configurations. Fig. 1 provides a graphical summary of the phase diagram discussed here.
Figure 1: Phase diagram of the power-law banded random matrix ensemble, showing the phases of mid-spectrum eigenstates as a function of the power-law exponent $\alpha$. The panels show the eigenstate entanglement entropy for a single realization as a function of the normalized eigenstate index (random labeling scheme, $L=10$), for the three phases ($\alpha=0.25$, $\alpha=0.75$, $\alpha=1.25$). In the panels, the ground state is on the left of the horizontal axis, while the highest excited eigenstate is on the right. The horizontal dashed line indicates the Page entanglement entropy $S_{\text{Page}}$. (Ensemble-averaged versions of such plots are discussed later, see Fig. 4.)
The phenomenology of power-law banded random matrices can be understood from resonance counting [102, 103] and the Breit-Wigner approximation (see, for example, Refs. [104, 21]). First, from resonance counting arguments, it follows that localization perturbation theory converges if the number $N_{\text{res}}$ of resonances, defined as $|H_{mn}|>|H_{nn}-H_{mm}|$, does not grow with the size $N$ of the matrix [23, 25, 65]. To calculate $N_{\text{res}}$, one should sum the probabilities of the other nodes $m\neq n$ to be in resonance with node $n$. For independent diagonal elements $H_{ii}$ sampled identically and uniformly from [W/2,W/2][-W/2,W/2], one obtains
$N_{\text{res}}=\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}\int_{-W}^{W}P(H_{mn})|H_{mn}|dH_{mn}.$ (4)
In the case of a normal distribution of $H_{ii}$ with $W$ related to the standard deviation, the above expression is the same in the leading order up to an unimportant prefactor. For power-law banded random matrices, the above integral gives $\int_{-W}^{W}P(H_{mn})|H_{mn}|dH_{mn}\sim R^{-\alpha}$ with R=|nm|R=|n-m| and leads to $N_{\text{res}}\sim\sum_{R}R^{-\alpha}\sim N^{1-\alpha}+\mathcal{O}(1)$, which converges for $\alpha>1$. Here and further, the summation runs over all admissible values of $R$, of which there are $\mathcal{O}(N)$. Therefore, for $\alpha>1$ one expects localized eigenstates.
@@
$\Gamma_{n}(E)=\frac{2\pi}{\hbar}\nu(E)\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}|H_{mn}|^{2},$ (5)
@@
with a certain normalization prefactor $A$ [104, 21]. Here, $\psi_{E}(n)$ is the overlap of an eigenstate having eigenenergy $E$ with the basis state $|n\rangle$. The broadening $\Gamma_{n}(E)$ in general depends both on the eigenvalue $E$ and on the location of the wave-function maximum $n$, where the time-dependent perturbation theory has been started. As soon as $\Gamma(E)\equiv\left\langle\Gamma_{n}(E)\right\rangle\gg\left\langle H_{nn}^{2}\right\rangle$ for all the energies $E$, all the wave-function coefficients are of the same order and, therefore, all the states should be ergodic. For power-law banded random matrices, one obtains the Breit-Wigner broadening to be $\Gamma_{n}(E)\sim\sum_{R}R^{-2\alpha}\sim N^{1-2\alpha}+\mathcal{O}(1)$. This diverges at $\alpha<\frac{1}{2}$ for all the energies. Therefore, for $\alpha<\frac{1}{2}$, one expects all the eigenstates to be ergodic (statistically equivalent to those of the GOE). Together, these two lines of arguments predict the ‘phase diagram’ shown in Fig. 1.
@@
In the quantum many-body interpretation, to be developed in the following sections, the two transitions appear to correspond to entanglement phase transitions [68]. As one increases $\alpha$, we start off with all eigenstates having volume law behavior for $\alpha<\frac{1}{2}$. As $\alpha$ crosses $\alpha=\frac{1}{2}$, the edge eigenstates change in nature and acquire area law scaling, and then when $\alpha$ crosses $\alpha=1$, the bulk eigenstates also turn to area law scaling. These notions will be confirmed through numerical scaling analyses later on, but we have provided a visualization of the different behaviors in the three panels in Fig. 1. The panels show single-realization scatter plots of $S_{\text{ent}}$ against eigenvalue index, using what we call the random labeling scheme (Sec. III).